OpenStudy (hba):

If x is the average of k and 10,and y is the average of k and 4,what is the average of x and y,in terms of k ?

OpenStudy (hba):

@hartnn I do not even know how to translate this

hartnn (hartnn):

that made my head spin....

OpenStudy (hba):

x=k+10/2

hartnn (hartnn):

k+10 = 2x k+4=2y (x+y)/2 = ?

hartnn (hartnn):

add first 2 equations and divide by 4

OpenStudy (hba):

first teach me how to translate it ?

hartnn (hartnn):

x is the average of k and 10 x= (k+10)/2

OpenStudy (hba):

Oh i got it :D

hartnn (hartnn):

good :)

OpenStudy (hba):

k+7/2 @hartnn >Right or wrong ?

hartnn (hartnn):

(2k+14) /4 = k/2+7/2

OpenStudy (hba):

haha it is the same lol Thanks :)

OpenStudy (hba):

6.7times10^-2=0.067 right or wrong ? @hartnn

hartnn (hartnn):

right.